### v1

# Using Large Language Models for Automatic Hint Ranking and Generation

Abstract
The use of Large Language Models (LLMs) has increased significantly recently, with individuals frequently interacting with chatbots to receive answers to a wide range of questions. In an era where information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills. This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers. We first introduce a manually constructed hint dataset, WikiHint, which includes 5,000 hints created for 1,000 questions. We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts. We assess the effectiveness of the hints with human participants who try to answer questions with and without the aid of hints. Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings. Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves hint quality, and (c) encoder-based models perform better than decoder-based models in hint ranking.
1 Introduction
Figure 1: Pipeline of WikiHint dataset generation. The numbers in arrows indicate the counts of output questions.
In recent years, question answering (QA) systems have risen in importance, giving users the opportunity to ask arbitrary questions and gain answers to them Mavi et al. (2024); Karpukhin et al. (2020); Zhao et al. (2021); Abdel-Nabi et al. (2023). The rapid development of Large Language Models (LLMs) Gemini Team et al. (2023); OpenAI et al. (2023); Dubey et al. (2024) has without doubt contributed to this, as well as to many other Natural Language Problems (NLPs) Qin et al. (2024). While the benefits of using LLMs for current information access are clear [^1] While the current LLMs have still considerable weaknesses such as hallucinations, once can assume that, in the future, these problems will be mitigated to large extent, judging from the speed of the recent technology advancement., there are some worries in relation to their potential effect on human development. One such concern relates to the potential weakening of important cognitive skills of users like thinking, reasoning, and remembering due to the expected widespread use of automatic question answering technologies, in particular, ones backed by powerful AI solutions  Heersmink (2024). Users who will rely mainly on the solutions presented by AI’s, might also be discouraged to practice and improve their reasoning abilities Alfredo et al. (2024). For example, Darvishi et al. (2024) demonstrate that students are more likely to depend on AI assistance instead of learning from it. Moreover, Jošt et al. (2024) examine the impact of LLMs as an automated problem-solving technology on education and learning outcomes, demonstrating that such systems can negatively affect the development of learning skills. Furthermore, psychological studies confirm the importance of obtaining an answer independently, enhancing user self-confidence and encouraging further learning Bandura (2013). Letting users come up with the correct answers by themselves should then also contribute to the positive psychological effect, potentially increasing their self-confidence and motivation for learning Usher and Pajares (2006).
While there is no simple remedy to the aforementioned problem, we would like to promote an approach that involves humans in the answer finding process, by providing them with hints rather than direct answers. Hints are meant to serve as subtle clues to guide potential users towards a correct answer, without revealing the solution Hume et al. (1996). This should engage human’s cognitive abilities, potentially leading to forming new pathways in brains based on the received hints and already possessed knowledge. Automatically generating hints for user questions could be used as an alternative for those who prefer to find the answers themselves; much like long-distance walking is an alternative to using a car for people who wish to take additional effort for the benefit of staying healthy Panter et al. (2018).
In this paper, we propose the first dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint which has been manually constructed and designed for both hint generation as well as evaluation. We next explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models. We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches. Finally, we assess the effectiveness of a novel evaluation method for hint ranking called HintRank and compare it with other automatic evaluation techniques. To sum up, we make the following contributions in this paper:
  • We release the first manually created dataset called WikiHint for HG task containing 5,000 hints and 1,000 questions.
  • We propose an automatic evaluation method for ranking hints called HintRank and compare with other evaluation methods.
  • We finetune and evaluate LLMs on WikiHint to assess the dataset quality and LLMs capabilities in hint generation and ranking.
  • We present several novel observations, including the findings of a positive correlation between hint convergence and helpfulness, an inverse correlation between their length and helpfulness, and the superiority of the answer-aware approach over the answer-agnostic approach. In general, our research can contribute to fostering research in explainable AI for education Khosravi et al. (2022).
2 Related Work
Automatic question answering (QA) Karpukhin et al. (2020); Zhao et al. (2021); Abdel-Nabi et al. (2023) and question generation (QG) Kurdi et al. (2020); Lu and Lu (2021); Zhang et al. (2021) have advanced quite much in the last years. These tasks have seen numerous different datasets Trischler et al. (2017); Rajpurkar et al. (2016); Zhang et al. (2020), and evaluation metrics Nema and Khapra (2018); Mavi et al. (2022) proposed. The research related to hint generation is however still scarce despite that hinting is a common mechanism used by humans for question answering, and that automatic hint generation could be regarded as the third missing task alongside the two established ones, QA and QG. The prior research focused mainly on generating hints for programming Price et al. (2019); Kochmar et al. (2022); Barnes and Stamper (2008); McBroom et al. (2021) and typically in the context of intelligent tutoring systems.
Automatic hint generation for factoid questions was first addressed by Jatowt et al. (2023). However, the authors neither released a dataset nor utilized LLMs, focusing instead on hints generated from selected Wikidata [^3] https://www.wikidata.org/ predicates. Moreover, their work only considered an answer-aware setting and did not explore the hint ranking task. Subsequently, Mozafari et al. (2024) released the first synthetic dataset for hint generation (HG) called TriviaHG, which was automatically generated using LLMs. However, the automatic generation increases the likelihood of false information within the dataset, particularly due to the hallucination phenomenon of LLMs. The authors also introduced the first automatic evaluation method, called Convergence, for assessing hint quality. However, this method requires substantial computational resources as it relies on LLMs for evaluation. Interested readers can also refer to the recent survey of Jangra et al. (2024) who discuss various types of hints and challenges associated with hint generation and evaluation.
To the best of our knowledge, WikiHint is the first manually curated dataset for the HG task, with questions and hints verified by humans. Our proposed automatic evaluation method is also the first for this task that does not rely on LLMs and is lightweight enough to be used locally.
3 WikiHint Dataset
The absence of high-quality, verified hint datasets poses a significant challenge given the demanding data requirements of LLMs for their effective training. In this section, we outline the process for constructing WikiHint dataset. Figure 1 provides an overview of the pipeline of the dataset generation process, which we explore in detail in the following sections.
3.1 Question Sampling Module
We incorporated AI-generated questions using ChatGPT OpenAI (2022) and questions from existing popular QA datasets such as SQuAD 2.0 Rajpurkar et al. (2018) and Natural Questions (NQ) Kwiatkowski et al. (2019). The following prompt was used for question generation:
Can you give me 10 questions where the answer is ANSWER? Please put them in a CSV file with answer=ANSWER and link=WIKIPEDIA_LINK where each question has an answer and link. Make sure to put the questions in quotation marks.
where WIKIPEDIA_LINK is the URL for the Wikipedia page corresponding to ANSWER. As mentioned above, we also selected questions from SQuAD 2.0 and NQ making sure that their answers had dedicated Wikipedia articles with sufficiently long content. Finally, we manually verified the correctness of all the selected questions discarding questions that are either too general or have incorrect answers. Table 6 in Appendix A compares the difficulty levels of questions based on their sources.
Figure 2: The HintRank method.
3.2 Hint Creation
The crowdsourcing platform Amazon Mechanical Turk [^4] https://www.mturk.com/ was then used to distribute the hint creation task among multiple workers. The instructions shown to crowdworkers asked them to create five hints for a question and its associated Wikipedia article. After generating the hints, the workers were asked to rank them on a scale from 1 to 5, with 1 being the most helpful in finding correct answers and 5 being the least helpful. Figure 9 in Appendix A displays the annotators’ interface for hint generation while Figures 10, 11, and 12 provide additional views of the summarized instructions, detailed instructions, and the provided examples, respectively, to further assist the crowdworkers. Each data submission was subsequently reviewed manually for quality and was either approved or rejected based on their assessment. Table 14 in Appendix A shows the detailed criteria used for the selection process. The most common reasons for rejection were hints that directly revealed the answers (answer leakage) and hints that were single words instead of complete sentences. Among the 2,788 submissions reviewed, 1,788 were rejected and 1,000 were accepted. We prepared several attributes to be included for each question, answer, and hint, as shown in Tables 7, 8, and 9 in Appendix A. We discuss some of them in Section 4. Finally, we divided the questions and hints into train and test subsets. The train subset includes 4,500 hints for 900 questions, while the test subset contains 500 hints for 100 questions. Tables 10, 11, 12, and 13 in Appendix A show few examples of hints taken from the WikiHint dataset.
Train Test
Number of hints 4,500 500
Number of questions 900 100
Avg. question length (words) 19.55 19.19
Avg. hint length (words) 17.77 18.32
Avg. #entities / question 1.2 1.44
Avg. #entities / hint 1.2 1.18
Table 1: Statistics of WikiHint dataset.
Rank 1 2 3 4 5
Average Length 16.99 17.67 18.02 18.14 18.3
Table 2: Average length of hints vs. their ranks.
Dataset Subset Relevance Readability Convergence Familiarity Length
Answer Leakage
Degree (Avg)
Answer Leakage
Degree (Max)
TriviaHG Entire 0.95 0.71 0.57 0.77 20.82 0.23 0.44
WikiHint Entire 0.98 0.72 0.73 0.75 17.82 0.24 0.49
TriviaHG Train 0.95 0.73 0.57 0.75 21.19 0.22 0.44
WikiHint Train 0.98 0.71 0.74 0.76 17.77 0.24 0.49
TriviaHG Test 0.95 0.73 0.6 0.77 20.97 0.23 0.44
WikiHint Test 0.98 0.83 0.72 0.73 18.32 0.24 0.47
Table 3: Quality comparison of WikiHint and TriviaHG. Relevance, convergence, familiarity, and answer leakage are measured on a scale from 0 to 1, while readability is rated on a scale from 0 to 2 (the lower, the more readable).
4 Evaluation Approaches
Mozafari et al. (2024) proposed several evaluation metrics for assessing the quality of hints, including Relevance, Readability, Convergence, and Familiarity, although the authors proposed automatic evaluation methods only for Convergence and Familiarity. Convergence is a measure of how effectively a hint can narrow down or eliminate potential answers to a given question. Familiarity measures the expected level of knowledge of information expressed in hints. To evaluate hints based on Convergence and Familiarity, we follow the method proposed by Mozafari et al. (2024). However, we employ two cores for convergence including LLaMA-3.1-8b and LLaMA-3.1-70b Dubey et al. (2024).
We extend the above evaluation scheme by incorporating automatic methods for evaluating hint’s Relevance and Readability. We also propose a new metric for evaluating the probability of answer leakage - a case when a hint directly reveals the answer in its content. We introduce a lightweight automatic evaluation method for assessing hint quality in a pairwise scenario. The way to compute those additional metrics is briefly described below.
Hints can be considered a form of an answer since they provide explanations of the question’s correct answer. Based on this, one can evaluate the Relevance of a hint to its question as an Answer Relevance task Es et al. (2024) - the task where the goal is to assess how pertinent the provided answer is to the target question. To compute the answer relevance metric, we employ DeepEval framework [^5] https://docs.confident-ai.com/ treating hint as a kind of answer. To evaluate Readability Liu and Lee (2023), we finetune a RoBERTa Liu et al. (2019) model as a classifier on the OneStopEnglish dataset Vajjala and Lučić (2018). The finetuned model categorizes sentences into three classes: Beginner (0), Intermediate (1), and Advanced (2), reflecting their level of reading difficulty [^6] The accuracy of the readability estimator model is 62.3%.
To calculate the Answer Leakage Degree, we measure the semantic similarity between each word of a hint and an answer using RoBERTa model.
4.1 HintRank Evaluation Approach
In addition to the above automatic evaluation approaches involving individual hints, we introduce a new lightweight evaluation method, HintRank, for evaluating and ranking hints using pairwise preferences. Building on the success of widely-used automatic evaluation metrics like BERTScore Zhang et al. (2020), BEM Bulian et al. (2022), MoverScore Zhao et al. (2019), and BLEURT Sellam et al. (2020), which leverage BERT Devlin et al. (2019) as the core evaluation module and demonstrate its effectiveness, we chose BERT as the foundation for the HintRank method. Our method determines the better hint within a pair of hints. Figure 2 illustrates the proposed method. In the HintRank method, we begin by concatenating a given question and its answer with two hints, labeled as $\text{Hint}_{1}$ and $\text{Hint}_{2}$, to create an input compatible with BERT model. Note that in the answer-agnostic scenarios, we avoid appending the answer to the evaluated hints. Such constructed input is then processed by BERT model, which produces one of two possible outputs: 0 or 1. An output of 0 means that $\text{Hint}_{2}$ is of higher quality than $\text{Hint}_{1}$, whereas an output of 1 suggests that $\text{Hint}_{1}$ is superior to $\text{Hint}_{2}$. As HintRank operates on pairwise preferences, it requires $\binom{n}{2}$ comparisons for a question with $n$ hints, with a runtime complexity of $\mathcal{O}(n^{2})$.
5 Experiments and Results
5.1 Data Analysis
The WikiHint dataset is split into a train set with 4,500 hints (900 questions) and a test set with 500 hints (100 questions). Table 1 provides the statistics of both train and test sets, while Figure 7 in Appendix A shows their distributions according to the question types, indicating that the distributions are well-matched.
Figure 3: Average convergence of the hints of WikiHint based on the hint ranks.
We next analyze the difficulty levels of questions in WikiHint. To evaluate the difficulty, we utilize the Reference-based Question Complexity method Gabburo et al. (2024). This method computes the difficulty of a question by assessing how many of its retrieved passages contain the correct answer and by measuring the relevance between the retrieved passages and the question. It then calculates the difficulty score for the question based on such computed features. In particular, we use the DPR method Karpukhin et al. (2020) as the retrieval technique, employing an English Wikipedia dump preprocessed by Karpukhin et al. (2020) as the evidence source, and consider the top 30 most relevant passages as the retrieved passages. Figure 8 in Appendix A illustrates the computed question difficulty of WikiHint for train and test subsets [^7] We classify questions with difficulty scores below 0.33 as easy, those above 0.66 as hard and the rest as medium.. The figure indicates that medium-hard questions are the most common as well as the train and test subsets have quite similar distributions in terms of question difficulty. Also, Table 6 in Appendix A highlights the difficulty levels of questions generated or extracted from various sources.
Table 2 reveals an interesting insight regarding the length of hints, which can be considered as one of indicators of helpfulness. The results suggest that high-quality hints tend to be shorter in length (measured by the number of words) than the lower quality hints. This finding indicates an inverse correlation between hint length and helpfulness, challenging the intuition that longer hints are more informative or specific, and therefore more useful. In contrast, shorter hints appear to be more concise and easier to follow, likely presenting more helpful information in the first place.
We also evaluate the hints in the entire WikiHint dataset (and separately, in its train and test subsets) using the relevance, readability, convergence, familiarity, length, and answer leakage degree. We then compare these values with the ones obtained for TriviaHG dataset Mozafari et al. (2024) - the only existing hint dataset. The comparison results are presented in Table 3. The results indicate that in terms of relevance, readability, answer leakage degree, and familiarity, the metrics are nearly same between the two datasets. However, WikiHint has better convergence values compared to TriviaHG. Additionally, the hints in WikiHint are shorter in length, as measured by word count, than TriviaHG. These results indicate that the hints in WikiHint are of higher quality.
Lastly, Figure 3 demonstrates the negative correlation between the convergence scores of hints and their helpfulness as represented by hint ranks assigned by crowdworkers. The plot suggests that the convergence scores can be considered a reliable metric for evaluating the helpfulness of hints and for hint ranking.
5.2 Human Evaluation
To manually evaluate hints, we recruited five independent evaluators, who were not involved in the dataset generation process, to answer the questions from the test subset of the WikiHint. The process was as follows: 1. Participants were asked to answer the question without using any hints. If they provided a correct answer, they proceeded to the next question. 2. If they could not answer the question correctly, they were asked to review the hints until they could find the correct answer. By providing the correct answer, the participants could move to the next question. 3. If the participants could not answer the question after reviewing all the hints, they were allowed to skip the question.
Figure 4: The results of human evaluation.
Model Config
Use
Answer
Relevance Readability
Convergence
(LLaMA 8b)
Convergence
(LLaMA 70b)
Familiarity Length
Answer Leakage
Degree (Avg)
Answer Leakage
Degree (Max)
GPT-4 Vanilla 0.91 1.0 0.14 0.48 0.84 26.36 0.23 0.51
GPT-4 Vanilla 0.92 1.1 0.12 0.47 0.81 26.93 0.24 0.52
LLaMA-3.1-405b Vanilla 0.94 1.49 0.11 0.47 0.76 41.81 0.23 0.5
LLaMA-3.1-405b Vanilla 0.92 1.53 0.1 0.45 0.78 50.91 0.23 0.5
LLaMA-3.1-70b FTwA 0.88 1.5 0.09 0.42 0.84 43.69 0.22 0.48
LLaMA-3.1-70b Vanilla 0.86 1.53 0.05 0.42 0.8 45.51 0.23 0.5
LLaMA-3.1-70b FTwoA 0.86 1.5 0.08 0.38 0.8 51.07 0.22 0.51
LLaMA-3.1-70b Vanilla 0.87 1.56 0.06 0.38 0.76 53.24 0.22 0.5
LLaMA-3.1-8b FTwA 0.78 1.63 0.05 0.37 0.79 50.33 0.22 0.52
LLaMA-3.1-8b Vanilla 0.81 1.72 0.05 0.32 0.8 54.38 0.22 0.5
LLaMA-3.1-8b FTwoA 0.76 1.7 0.03 0.32 0.8 55.02 0.22 0.51
LLaMA-3.1-8b Vanilla 0.78 1.76 0.04 0.3 0.83 52.99 0.22 0.5
Table 4: Evaluation of generated hints based on relevance, readability, convergence, familiarity, length, and answer leakage across different scenarios.
Figure 4 illustrates that all the participants could answer more questions across all of question types such as HUMAN, ENTITY, and LOCATION [^8] We use names as stated in the original dataset. when they used hints compared to the case without hints. Notably, the greatest improvement was observed in human-related questions, where hints proved most beneficial. Following, entity-related questions led to significant improvement, while location-related questions saw the smallest positive change. This suggests that generating effective hints becomes progressively more challenging for human, entity, and location questions, in that order.
5.3 Model Performance
To further assess the quality of hints, we analyze how well LLMs can automatically generate hints for questions. We use the open-source LLaMA models: LLaMA-3.1-8b, LLaMA-3.1-70b, and LLaMA-3.1-405b Dubey et al. (2024), and GPT-4 OpenAI et al. (2023) as the most powerful closed-source LLM for comparison.
To explore different scenarios, we finetune [^9] We perform model finetuning using the API functions available on together.ai LLaMA-3.1-8b and LLaMA-3.1-70b on the training subset of the WikiHint dataset to evaluate the LLMs’ capabilities in hint generation when trained specifically on this task. For each question, we assign a hint as the target during the finetuning process. As a result of this learning strategy, during the inference stage, the finetuned model is prompted to generate one hint for each question. We consider two finetuning approaches: answer-aware and answer-agnostic. Given that LLMs typically handle most knowledge questions correctly, the answer-agnostic approach might be sufficient for generating hints. Besides, users generally do not know the answers to their questions when seeking hints. However, the answer-aware approach has its own advantages, too, such as in educational contexts where a teacher might use it to collect materials for class preparation. Due to the importance of both approaches, we chose to investigate fine-tuning of the LLMs in these two distinct scenarios.
We found that shorter prompts were more effective in achieving the desired task. Longer, more detailed instructions often led to the model disregarding the key goal, i.e., generating hints, and instead focusing on irrelevant details. In contrast, shorter prompts increased the likelihood of successful task completion. After experimenting, we opted for the following prompt as the system prompt:
You are a hint generator for the factoid questions. The user asks you a question and you should generate a hint for that question without revealing the answer in the hint.
Two distinct user prompts were employed to generate hints within a zero-shot learning strategy. Assuming a question $q$ as an input, the answer-agnostic prompt was ‘Give me the best hint for this question: $q$. The answer-aware prompt included the answer $a$ as follows: ‘Give me the best hint for this question: $q$? The answer for the question is $a$’.
Figure 5: Average length of generated hints by LLMs.
To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA where wA means With Answer and woA means Without Answer. We test these models on the WikiHint test to assess the impact of finetuning and the inclusion of answers in the prompt.
Figure 5 illustrates that as LLMs decrease in size and hint generation capability, the length of the generated hints increases. This supports our observation made in Section 5.1 of an inverse correlation between hint length and hint quality. Additionally, hints produced by finetuned models are generally shorter than those from vanilla models, indicating that finetuned models may generate higher-quality hints. Moreover, hints in the answer-aware scenarios are shorter compared to those in answer-agnostic scenarios, suggesting that when the answer is provided along with the question, LLMs are able to produce more effective hints.
Table 4 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity. The results indicate that more powerful LLMs are capable of generating more relevant hints. Regarding readability, GPT-4 exhibits the highest quality, followed closely by LLaMA-3.1-405b and LLaMA-3.1-70b, while LLaMA-3.1-8b shows the lowest readability. It also demonstrates that more powerful LLMs can generate more readable hints. Additionally, finetuned models consistently outperform their vanilla counterparts, and answer-aware prompts yield better results compared to answer-agnostic prompts for readability and familiarity. The answer leakage degree indicates that the prompt we use is effective in preventing LLMs from including answers, their synonyms or very similar terms in the generated hints, as the results closely align with those of WikiHint shown in Table 3.
Figure 6 illustrates that as LLMs decrease in size and capability, the convergence of their hints also diminishes. This trend is observed for both LLaMA-3.1-8b and LLaMA-3.1-70b, used as the cores of the convergence method. This supports our claim in Section 5.1 regarding the correlation between convergence and ranks. The figure also shows that the average convergence for the answer-aware approach surpasses that of the answer-agnostic approach, suggesting that including the answer in the prompt makes it easier for LLMs to generate hints. Furthermore, LLMs finetuned on the train subset of the WikiHint dataset achieve better convergence scores than their vanilla counterparts, indicating the efficacy of WikiHint for finetuning LLMs in hint generation.
Figure 6: Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
5.4 HintRank Evaluation Method
Method Config Use Answer Accuracy (%) Correlation (%)
Convergence Vanilla 40.80 36.70
LLaMA-3.1-8b Vanilla 60.50 49.25
LLaMA-3.1-8b Vanilla 60.95 49.79
LLaMA-3.1-8b FTwoA 61.00 50.74
LLaMA-3.1-8b FTwA 61.25 49.03
LLaMA-3.1-70b Vanilla 64.00 50.32
LLaMA-3.1-70b Vanilla 64.25 51.32
LLaMA-3.1-70b FTwoA 64.65 51.51
LLaMA-3.1-70b FTwA 65.30 52.53
HintRank FTwoA 67.25 49.06
HintRank FTwA 68.55 52.34
Table 5: Comparison between Convergence metric, LLM-based ranking, and HintRank.
As outlined in Section 4, we also propose in this paper a novel evaluation method, HintRank, for ranking hints using the BERT model. Alongside finetuning BERT, we additionally finetune LLaMA-3.1-8b and LLaMA-3.1-70b models on the train set of the WikiHint to assess the performance of these LLMs in identifying high-quality hints. Similar to the experiments described in Section 5.3, we examine various scenarios including answer-aware and answer-agnostic contexts, and compare vanilla models with their finetuned counterparts. We use the following prompt as the system prompt:
You are a hint evaluator for the factoid questions. The user gives you a question and two hints and you should specify which hint for that question is a better hint and more helpful.
Two distinct user prompts are employed to evaluate hints within a zero-shot learning strategy. Assuming a question $q$ as a question and $h_{1}$ and $h_{2}$ as a pair of hints, the answer-agnostic prompt is:
Which hint is better to find the answer of this question: $q$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
and the answer-aware prompt with answer $a$ is:
Which hint is better to find the answer of this question: $q$. The answer for this question is $a$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
We benchmark HintRank against the Convergence metric which turned out to be useful for hint ranking assessment as indicated in Figure 3. To convert pairwise rankings to listwise rankings, we apply the Bradley–Terry model Bradley and Terry (1952). We evaluate the correlation between the rankings with Pearson Correlation Mining (2006).
Table 5 outlines the key features and differences among various scenarios. The results indicate that with the increase in the size and power of LLMs, both accuracy and correlation improve. Additionally, the answer-aware approach yields better outcomes compared to the answer-agnostic method, suggesting that the presence of an answer enables LLMs to evaluate hints more effectively. Moreover, finetuned versions outperform their vanilla counterparts, demonstrating that the WikiHint dataset is well-suited for model fine-tuning to rank hints. Surprisingly, the BERT-base method outperforms LLMs, including finetuned versions in the answer-aware scenario. This holds true for both Bert-FTwoA and Bert-FTwA, although BERT-base also performs better in the answer-aware approach compared to answer-agnostic. BERT-base methods achieve higher accuracy than LLMs and convergence, but in terms of correlation, LLaMA-3.1-70b exhibits the best performance. The effectiveness of BERT-base methods may be attributed to the strengths of encoder-based models like BERT in classification tasks over decoder-based models. Utilizing BERT-based models instead of LLMs enhances the speed and accessibility of HintRank, reducing computational demands. Figure 13 in Appendix A shows that accuracy improves as the rank difference between hints increases, indicating it’s harder to correctly order hints with closer ranks.
6 Conclusions
In this paper, we introduced the first manually created dataset for hint generation and hint ranking. We also presented a new lightweight method for evaluating and ranking hints. To demonstrate the effectiveness of our dataset, we conducted experiments where humans attempted to answer questions with and without the use of hints. The results confirm that the hints are of sufficient quality to assist users. We then finetuned LLMs using our dataset, prompting them to generate new hints for different questions. The high quality of the generated hints indicates that our dataset is well-suited for finetuning LLMs for HG task. We also finetuned BERT and LLMs on the dataset for the task of hint ranking and evaluated their performance. The results reveal that encoder-based models outperform decoder-based models in hint ranking.
In future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.
Limitations
Our study has the following limitations:
  • The need for generative capabilities in hint generation task necessitates the use of LLMs. However, this dependency is a limitation as fine-tuning and prompting LLMs require extensive computational resources and are time-consuming.
  • Our research focus on factoid questions may limit its applicability to other types of questions that involve more complex or abstract answers. Factoid questions, by their nature, provide clear and concrete answers, which simplifies automated hint generation and evaluation but may not fully capture the breadth of human inquiry.
  • The WikiHint dataset is exclusively written in the English language. While this facilitates accessibility for a global audience and ensures compatibility with most existing Large Language Models, it also limits the dataset’s applicability in multilingual or non-English contexts, potentially excluding non-English speakers and diverse linguistic data.
Ethical Considerations
Our study utilizes GPT models, which are covered by the OpenAI License and Apache-2.0 license, and the LLaMA model, which is distributed under Meta’s LLaMA 2 Community License Agreement. We comply with these licensing agreements in all applications. Additionally, the datasets we use are sourced from repositories that are approved for academic use. The artifacts developed during our research are made available under the MIT license to facilitate straightforward modifications and use by the research community. We ensure that our data management, model training, and dissemination practices meet ethical standards and legal requirements associated with each artifact we use.
Appendix A Appendix
Figure 7: The distribution of Train and Test subsets.
Figure 8: Question difficulty based on different question types.
Source ChatGPT NQ SQuAD 2.0
Difficulty 0.43 0.34 0.38
Table 6: Question distributions based on the their sources and difficulty.
Figure 9: The MTurk Worker interface for the hint generation task.
Figure 10: The summarized instructions for the hint generation task.
Figure 11: The detailed instructions for the hint generation task.
Figure 12: Good and bad examples for the hint generation task.
Attribute Description
question The content of the question.
major The major category of the question.
minor The specific sub-category of the question.
entity Content of an entity
ent_type Type of entity (e.g., GPE, PERSON).
start_index Start index of the entity in the question.
end_index End index of the entity in the question.
wikipedia_page_title The title of the corresponding Wikipedia page to the entity in question.
wiki_views_per_month Number of views per month of the Wikipedia page for the entity in question.
normalized_views Views normalized to scale from 0 to 1 for the entity in the question.
readability Indicates the readability score
familiarity Indicates the familiarity score
difficulty Indicates the difficulty level of the question
Table 7: A detailed description of attributes of a question in WikiHint.
Attribute Description
answer The actual answer.
entity Content of an entity as identified within the answer.
ent_type Type of entity.
start_index Start index of the entity in the answer.
end_index End index of the entity in the answer.
wikipedia_page_title The title of the corresponding Wikipedia page to the entity in the answer.
wiki_views_per_month Number of views per month of the Wikipedia page for the entity in the answer.
normalized_views Views normalized to scale from 0 to 1 for the entity in the answer.
familiarity Indicates the familiarity score
difficulty Indicates the difficulty level of the answer
Table 8: A detailed description of attributes of an answer in WikiHint.
Attribute Description
hint Hint provided to assist with the question.
source URL source of the hint.
entity Content of entities mentioned in the hint.
ent_type Category of each entity (e.g., PERSON, GPE) mentioned in the hint.
start_index Specific start index where the entity is found in the hint text.
end_index Specific end index where the entity is found in the hint text.
wikipedia_page_title The title of the corresponding Wikipedia page to the entities in the hint.
wiki_views_per_month Number of views per month of the Wikipedia pages for the entities of the hint.
normalized_views Views normalized to scale from 0 to 1 for the entity in the hint.
relevance Indicates the relevance score
readability Indicates the readability score
convergence Indicates the convergence score
familiarity Indicates the familiarity score
answer_leakage Indicates the answer leakage score
rank Priority or helpfulness rating of the hint
Table 9: A detailed description of attributes of a hint in WikiHint.
Question What is the driest country in sub-Saharan Africa?
Fifth Best Hint The earliest settlers in this country were in the 18th century and crossed the Orange River to move into the area.
Fourth Best Hint The name of this country is derived from the oldest desert on Earth.
Third Best Hint This country’s coat of arms is a shield with the same design as the flag, 2 antelopes and a red blue and white bird.
Second Best Hint This country is a country in southern Africa with its western border along the Atlantic Ocean.
Best Hint This country’s flag is blue and green with a red stripe down the middle and the symbol of a sun in the top left corner.
Table 10: First example of the WikiHint.
Question What artist received the Presidential Medal of Freedom posthumously in 2018, in recognition of his contributions to American culture?
Fifth Best Hint The artist had karate as a lifelong interest, including its moves in his performances.
Fourth Best Hint The artist served in the Army after he became famous.
Third Best Hint The artist never performed outside North America.
Second Best Hint The artist tried to establish a career in films with pictures such as Jailhouse Rock and Fun in Acapulco.
Best Hint The artist is known for hits such as "Love Me Tender" and "Jailhouse Rock".
Table 11: Second example of the WikiHint.
Question What British comedy team is famous for its ’Four Yorkshiremen’ sketch, portraying exaggerated tales of hardship with humor?
Fifth Best Hint A giant cupid’s foot is repeatedly used in the show.
Fourth Best Hint This British comedy troupe is formed in 1969 consisting of 6 members.
Third Best Hint The team was awarded the AFI Star Award by the American Film Institute in 1998.
Second Best Hint The team received the BAFTA Award for Outstanding British Contribution to Cinema at the 41st British Academy Film Awards in 1988.
Best Hint The Holy Grail and Life of Brian are some of their greatest comedy films.
Table 12: Third example of the WikiHint.
Question What fictional pirate is known for carrying a compass that doesn’t point north but rather to what the user wants most?
Fifth Best Hint One of the pirate’s most famous quotes is "Now, bring me that horizon".
Fourth Best Hint The fictional pirate appeared in the video game series "Kingdom Hearts".
Third Best Hint The fictional pirate has a blood debt to another character called Davy Jones.
Second Best Hint The ship of the fictional pirate is called the Black Pearl.
Best Hint The fictional pirate is played by Johnny Depp in the film series.
Table 13: Fourth example of the WikiHint.
A hint must not include the exact answer explicitly. A hint must be a sentence. A hint must be specific, not generic. A hint must be from the corresponding Wikipedia page. A hint must have a unique rank.
Table 14: A detailed criteria or standards used for verifying the generated hints during the selection process.
Figure 13: Accuracy of HintRank for different hint pairs in different scenarios. The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.

### v2

@@
-Using Large Language Models for Automatic Hint Ranking and Generation
-Abstract
-The use of Large Language Models (LLMs) has increased significantly recently, with individuals frequently interacting with chatbots to receive answers to a wide range of questions.
-In an era where information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills.
+WikiHint: A Human-Annotated Dataset for Hint Ranking and Generation
+Abstract.
+The use of Large Language Models (LLMs) has increased significantly with users frequently asking questions to chatbots.
+In the time when information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills.
 This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers.
-We first introduce a manually constructed hint dataset, WikiHint, which includes 5,000 hints created for 1,000 questions.
+We first introduce a manually constructed hint dataset, WikiHint, which is based on Wikipedia and includes 5,000 hints created for 1,000 questions.
 We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts.
-We assess the effectiveness of the hints with human participants who try to answer questions with and without the aid of hints.
+We assess the effectiveness of the hints with human participants who answer questions with and without the aid of hints.
 Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings.
-Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves hint quality, and (c) encoder-based models perform better than decoder-based models in hint ranking.
-1 Introduction
-Figure 1: Pipeline of WikiHint dataset generation. The numbers in arrows indicate the counts of output questions.
-In recent years, question answering (QA) systems have risen in importance, giving users the opportunity to ask arbitrary questions and gain answers to them Mavi et al.
-(2024); Karpukhin et al.
-(2020); Zhao et al.
-(2021); Abdel-Nabi et al.
-(2023).
-The rapid development of Large Language Models (LLMs) Gemini Team et al.
-(2023); OpenAI et al.
-(2023); Dubey et al.
-(2024) has without doubt contributed to this, as well as to many other Natural Language Problems (NLPs) Qin et al.
-(2024).
-While the benefits of using LLMs for current information access are clear [^1] While the current LLMs have still considerable weaknesses such as hallucinations, once can assume that, in the future, these problems will be mitigated to large extent, judging from the speed of the recent technology advancement., there are some worries in relation to their potential effect on human development.
-One such concern relates to the potential weakening of important cognitive skills of users like thinking, reasoning, and remembering due to the expected widespread use of automatic question answering technologies, in particular, ones backed by powerful AI solutions  Heersmink (2024).
-Users who will rely mainly on the solutions presented by AI’s, might also be discouraged to practice and improve their reasoning abilities Alfredo et al.
-(2024).
+Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves quality of generated hints, and (c) encoder-based models perform better than decoder-based models in hint ranking.
+Hint Dataset, Hint Evaluation, Hint Generation
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
+1. Introduction
+Figure 1.
+Hints for a question from the WikiHint dataset, with their corresponding rankings (1 being the highest and 5 the lowest) which let users find the answer.
+Note that the opposite arrangement of the hints (from 1 to 5) would make the answer finding task easiest.
+Figure 2. Pipeline of WikiHint dataset generation. The numbers in the arrows indicate the counts of output questions. Qs and Hs denote the number of questions and hints, respectively.
+In recent years, question answering (QA) systems have risen in importance, offering users the opportunity to effortlessly receive answers to different types of questions (Mavi et al., 2024; Karpukhin et al., 2020; Zhao et al., 2021; Abdel-Nabi et al., 2023).
+The rapid development of Large Language Models (LLMs) (Gemini Team et al., 2023; OpenAI et al., 2023; Dubey et al., 2024) has mainly contributed to this, impacting also many other Natural Language tasks (NLPs) (Qin et al., 2024).
+While the benefits of using LLMs for current information access are quite evident, concerns have been raised in relation to their potential effect on human cognitive development.
+One such worry relates to the potential weakening of important cognitive skills of users like thinking, reasoning, and remembering due to the expected widespread use of the automatic question answering technologies, in particular, ones backed by powerful AI technologies  (Heersmink, 2024).
+Users who will rely mainly on the solutions presented by AI’s might also be discouraged to practice and improve their reasoning abilities (Alfredo et al., 2024).
 For example, Darvishi et al.
 (2024) demonstrate that students are more likely to depend on AI assistance instead of learning from it.
-Moreover, Jošt et al.
+Jošt et al.
 (2024) examine the impact of LLMs as an automated problem-solving technology on education and learning outcomes, demonstrating that such systems can negatively affect the development of learning skills.
-Furthermore, psychological studies confirm the importance of obtaining an answer independently, enhancing user self-confidence and encouraging further learning Bandura (2013).
-Letting users come up with the correct answers by themselves should then also contribute to the positive psychological effect, potentially increasing their self-confidence and motivation for learning Usher and Pajares (2006).
-While there is no simple remedy to the aforementioned problem, we would like to promote an approach that involves humans in the answer finding process, by providing them with hints rather than direct answers.
-Hints are meant to serve as subtle clues to guide potential users towards a correct answer, without revealing the solution Hume et al.
-(1996).
-This should engage human’s cognitive abilities, potentially leading to forming new pathways in brains based on the received hints and already possessed knowledge.
-Automatically generating hints for user questions could be used as an alternative for those who prefer to find the answers themselves; much like long-distance walking is an alternative to using a car for people who wish to take additional effort for the benefit of staying healthy Panter et al.
-(2018).
-In this paper, we propose the first dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint which has been manually constructed and designed for both hint generation as well as evaluation.
-We next explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models.
+Furthermore, psychological studies confirm the importance of obtaining an answer independently, enhancing user self-confidence (Bandura, 2013).
+Letting users come up with the correct answers by themselves should then also contribute to the positive psychological effect, potentially increasing their self-confidence (Usher and Pajares, 2006).
+While the problem of over-reliance on chatbots is complex and still not researched well, we wish to promote approaches that involve humans in the answer finding process.
+Providing users with hints rather than direct answers is one such approach.
+Hints are designed to offer subtle clues that guide users toward the correct answer without explicitly revealing it (Hume et al., 1996).
+Unlike the concept of framing (aka ”scattered conceptualization”) (Entman, 1993) which elevates particular pieces of information to promote a certain view or cause bias (Garrido-Muñoz et al., 2021; Färber et al., 2020), hints help humans to recall previously known information or guide them in the reasoning process.
+In this study, we define hints for factoid questions [^1] Factoid questions are those with answers in the form of entities such as locations, persons, etc.
+as clues that encourage users to think critically, reason, and draw on their knowledge to deduce the answer rather than having it explicitly provided.
+This approach aims to support and enhance cognitive skills such as critical thinking, reasoning, and memory.
+To ensure that hints are meaningful and widely accessible, we generate them based on external knowledge and well-established facts.
+For this purpose, we select Wikipedia as the primary source, as it is one of the most widely recognized and comprehensive repositories of factual information.
+In this paper, we propose the first manually curated dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint, which has been constructed and designed for hint generation, ranking and evaluation.
+Figure 1 shows a question from the WikiHint dataset, with its hints and the corresponding ranks reflecting the usefulness of each hint.
+We explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models.
 We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches.
-Finally, we assess the effectiveness of a novel evaluation method for hint ranking called HintRank and compare it with other automatic evaluation techniques.
-To sum up, we make the following contributions in this paper:
-  • We release the first manually created dataset called WikiHint for HG task containing 5,000 hints and 1,000 questions.
+Finally, we assess the effectiveness of a novel evaluation method for hint ranking, called HintRank, and compare it with other automatic evaluation techniques.
+Ranking hints allows arranging them either from easiest to hardest which can be used for selecting the most useful hints for a given question, or from hardest to easiest (as shown in Fig.
+1).
+The latter arrangement could find applications in quizzes, knowledge assessment scenarios, or in certain entertainment use cases (e.g., in games).
+In summary, we make the following contributions in this paper:
+  • We release the first manually created dataset called WikiHint for the HG task containing 5,000 hints and 1,000 questions.
   • We propose an automatic evaluation method for ranking hints called HintRank and compare with other evaluation methods.
-  • We finetune and evaluate LLMs on WikiHint to assess the dataset quality and LLMs capabilities in hint generation and ranking.
-  • We present several novel observations, including the findings of a positive correlation between hint convergence and helpfulness, an inverse correlation between their length and helpfulness, and the superiority of the answer-aware approach over the answer-agnostic approach.
-In general, our research can contribute to fostering research in explainable AI for education Khosravi et al.
-(2022).
-2 Related Work
-Automatic question answering (QA) Karpukhin et al.
-(2020); Zhao et al.
-(2021); Abdel-Nabi et al.
-(2023) and question generation (QG) Kurdi et al.
-(2020); Lu and Lu (2021); Zhang et al.
-(2021) have advanced quite much in the last years.
-These tasks have seen numerous different datasets Trischler et al.
-(2017); Rajpurkar et al.
-(2016); Zhang et al.
-(2020), and evaluation metrics Nema and Khapra (2018); Mavi et al.
-(2022) proposed.
-The research related to hint generation is however still scarce despite that hinting is a common mechanism used by humans for question answering, and that automatic hint generation could be regarded as the third missing task alongside the two established ones, QA and QG.
-The prior research focused mainly on generating hints for programming Price et al.
-(2019); Kochmar et al.
-(2022); Barnes and Stamper (2008); McBroom et al.
-(2021) and typically in the context of intelligent tutoring systems.
+  • We finetune and evaluate diverse LLMs on WikiHint to assess the dataset quality and LLMs’ capabilities in hint generation and ranking.
+  • We present several novel observations, including the findings of a positive correlation between hint convergence and helpfulness, an inverse correlation between hint length and helpfulness, and the superiority of the answer-aware approach over the answer-agnostic approach in hint generation.
+2. Related Work
+Automatic question answering (QA) (Karpukhin et al., 2020; Zhao et al., 2021; Abdel-Nabi et al., 2023) and question generation (QG) (Kurdi et al., 2020; Lu and Lu, 2021; Zhang et al., 2021) have progressed a lot in the last years.
+These tasks have seen numerous different datasets (Trischler et al., 2017; Rajpurkar et al., 2016; Zhang et al., 2020), and evaluation metrics (Nema and Khapra, 2018; Mavi et al., 2022) proposed.
+The research related to hint generation is however still scarce despite the fact that hinting is a common mechanism used by humans for question answering, and that automatic hint generation could be regarded as the third missing task alongside QA and QG.
+The prior research however focused mainly on generating hints for programming (Price et al., 2019; Kochmar et al., 2022a; Barnes and Stamper, 2008; McBroom et al., 2021).
+In the context of intelligent tutoring systems, automated, machine-learning-based methods have been shown to be effective in generating feedback and hints, reducing reliance on expert-crafted rules (Kochmar et al., 2020, 2022b).
+However, the effectiveness of LLMs in pedagogical tasks remains an open challenge.
+For instance, Tack and Piech (2022) investigated the teaching capabilities of LLMs such as Blender (Jiang et al., 2023) and GPT-3 (Brown et al., 2020) in educational dialogues and found that, while these models perform well in conversational uptake, they lack pedagogical effectiveness, particularly in providing helpful feedback.
+To address this limitation, several studies have proposed different approaches to improving LLM-generated feedback.
+For example, Kulshreshtha et al.
+(2022) leveraged a few-shot question generation strategy to create personalized feedback for mathematics problems, demonstrating improved student learning outcomes.
+Similarly, McNichols et al.
+(2023) explored the use of in-context learning and various LLMs for distractor generation in multiple-choice mathematics questions but found that the models still exhibit weaknesses, requiring further research for improvement.
+Additionally, Tonga et al.
+(2024) investigated the impact of different LLMs and prompting strategies on the quality of generated feedback and hints, highlighting that the effectiveness of LLM-generated assistance is highly dependent on the choice of model and the prompting approach used.
 Automatic hint generation for factoid questions was first addressed by Jatowt et al.
@@
 However, the authors neither released a dataset nor utilized LLMs, focusing instead on hints generated from selected Wikidata [^3] https://www.wikidata.org/ predicates.
-Moreover, their work only considered an answer-aware setting and did not explore the hint ranking task.
+Moreover, their work only considered an answer-aware setting in hint generation and did not explore the hint ranking task.
 Subsequently, Mozafari et al.
-(2024) released the first synthetic dataset for hint generation (HG) called TriviaHG, which was automatically generated using LLMs.
+(2024b) released the first synthetic dataset for hint generation (HG) called TriviaHG, which was automatically generated using LLMs.
 However, the automatic generation increases the likelihood of false information within the dataset, particularly due to the hallucination phenomenon of LLMs.
 The authors also introduced the first automatic evaluation method, called Convergence, for assessing hint quality.
-However, this method requires substantial computational resources as it relies on LLMs for evaluation.
+However, that method requires substantial computational resources as it relies on LLMs for evaluation.
+Mozafari et al.
+(2024a) also demonstrated that hints automatically generated from TriviaHG can serve as a context for QA systems, outperforming alternative approaches such as retrieval-based and generative methods in RAG.
 Interested readers can also refer to the recent survey of Jangra et al.
 (2024) who discuss various types of hints and challenges associated with hint generation and evaluation.
-To the best of our knowledge, WikiHint is the first manually curated dataset for the HG task, with questions and hints verified by humans.
-Our proposed automatic evaluation method is also the first for this task that does not rely on LLMs and is lightweight enough to be used locally.
-3 WikiHint Dataset
-The absence of high-quality, verified hint datasets poses a significant challenge given the demanding data requirements of LLMs for their effective training.
+To our knowledge, WikiHint is the first data set for the HG task with questions and their hints verified by humans.
+Our proposed automatic evaluation method for hint ranking is also the first for this task that does not rely on LLMs and is lightweight enough to be used locally.
+3. WikiHint Dataset
+The absence of verified high-quality hint datasets poses a significant challenge given the demanding data requirements of LLMs for their effective training.
 In this section, we outline the process for constructing WikiHint dataset.
 Figure 2 provides an overview of the pipeline of the dataset generation process, which we explore in detail in the following sections.
-3.1 Question Sampling Module
-We incorporated AI-generated questions using ChatGPT OpenAI (2022) and questions from existing popular QA datasets such as SQuAD 2.0 Rajpurkar et al.
-(2018) and Natural Questions (NQ) Kwiatkowski et al.
-(2019).
-The following prompt was used for question generation:
+Figure 3.
+The MTurk Worker interface for generating and ranking hints.
+\${question} represents the question presented to the worker, \${answer} is the corresponding answer, and \${link} provides the link to the relevant Wikipedia page.
+3.1. Question Sampling Module
+We first select questions that meet our criteria for hint generation.
+3.1.1. Question Selection
+We incorporated 4,000 questions, including randomly selected questions from popular QA datasets such as SQuAD 2.0 (Rajpurkar et al., 2018) and Natural Questions (NQ) (Kwiatkowski et al., 2019), as well as AI-generated questions using ChatGPT (OpenAI, 2022).
+The following prompt was used for question generation with ChatGPT:
 Can you give me 10 questions where the answer is ANSWER?
@@
 Make sure to put the questions in quotation marks.
-where WIKIPEDIA_LINK is the URL for the Wikipedia page corresponding to ANSWER.
+Figure 4. The instructions for the hint generation and ranking tasks on the Amazon MTurk platform.
+Table 1. A detailed description of attributes of a question.
+Attribute Description
+question The content of the question.
+major The major category of the question.
+minor The specific sub-category of the question.
+entity Content of an entity.
+ent_type Type of entity (e.g., GPE, PERSON).
+start_index Start index of the entity in the question.
+end_index End index of the entity in the question.
+wikipedia_page_title The title of the Wikipedia page corresponding to the entity in question.
+wiki_views_per_month Number of views per month of the Wikipedia page for the entity in question.
+normalized_views Views normalized to scale from 0 to 1 for the entity in the question.
+readability The readability score.
+familiarity The familiarity score.
+difficulty The difficulty level of the question.
+where WIKIPEDIA_LINK is the URL of the Wikipedia page corresponding to ANSWER.
 As mentioned above, we also selected questions from SQuAD 2.0 and NQ making sure that their answers had dedicated Wikipedia articles with sufficiently long content.
-Finally, we manually verified the correctness of all the selected questions discarding questions that are either too general or have incorrect answers.
-Table 6 in Appendix A compares the difficulty levels of questions based on their sources.
-Figure 2: The HintRank method.
-3.2 Hint Creation
+After this filtering process, 3,247 questions remained for the next steps.
+3.1.2. Question Verification
+We manually verified the correctness of all the selected questions, discarding those that were too general or contained incorrect answers.
+Additionally, we ensured that all questions included a question mark, adding it where necessary.
+We also reviewed all Wikipedia links and corrected any character errors in the URLs.
+After this filtering process, 2,788 questions remained for the next steps.
+3.2. Hint Generation Module
+In this module, we generate hints and rank them.
+3.2.1. Amazon MTurk
 The crowdsourcing platform Amazon Mechanical Turk [^4] https://www.mturk.com/ was then used to distribute the hint creation task among multiple workers.
 The instructions shown to crowdworkers asked them to create five hints for a question and its associated Wikipedia article.
-After generating the hints, the workers were asked to rank them on a scale from 1 to 5, with 1 being the most helpful in finding correct answers and 5 being the least helpful.
-Figure 9 in Appendix A displays the annotators’ interface for hint generation while Figures 10, 11, and 12 provide additional views of the summarized instructions, detailed instructions, and the provided examples, respectively, to further assist the crowdworkers.
-Each data submission was subsequently reviewed manually for quality and was either approved or rejected based on their assessment.
-Table 14 in Appendix A shows the detailed criteria used for the selection process.
+After generating the hints, the same workers were asked to rank them on a scale from 1 to 5, with 1 being the most helpful in finding correct answers and 5 being the least helpful.
+Figure 3 displays the annotators’ interface for hint generation while Figure 4 provides detailed instructions to further assist the crowdworkers.
+Table 2. A detailed description of attributes of an answer.
+Attribute Description
+answer The actual answer.
+entity Content of an entity as identified within the answer.
+ent_type Type of entity.
+start_index Start index of the entity in the answer.
+end_index End index of the entity in the answer.
+wikipedia_page_title The title of the Wikipedia page corresponding to the entity in the answer.
+wiki_views_per_month Number of views per month of the Wikipedia page for the entity in the answer.
+normalized_views Views normalized to scale from 0 to 1 for the entity in the answer.
+familiarity The familiarity score.
+difficulty The difficulty level of the answer.
+Table 3. A detailed description of attributes of a hint.
+Attribute Description
+hint Hint provided for the question.
+source URL source of the hint.
+entity Entities mentioned in the hint.
+ent_type Category of each entity (e.g., PERSON, GPE) mentioned in the hint.
+start_index Specific start index where the entity is found in the hint text.
+end_index Specific end index where the entity is found in the hint text.
+wikipedia_page_title The title of the Wikipedia page corresponding to the entities in the hint.
+wiki_views_per_month Number of views per month of the Wikipedia pages for the entities of the hint.
+normalized_views Views normalized to scale from 0 to 1 for the entity in the hint.
+relevance The relevance score.
+readability The readability score.
+convergence The convergence score.
+familiarity The familiarity score.
+answer_leakage The answer leakage score.
+rank Priority or helpfulness rating of the hint.
+3.2.2. Hint Verification
+Each data submission was then manually reviewed by three independent judges for quality and either approved or rejected based on their evaluation.
+The following section outlines the detailed criteria used in the selection process.
+  • The hint must not explicitly include the exact answer.
+  • The hint must be a sentence.
+  • The hint must be specific, not generic.
+  • The hint must be from the corresponding Wikipedia page.
+  • The hint must have a unique rank.
 The most common reasons for rejection were hints that directly revealed the answers (answer leakage) and hints that were single words instead of complete sentences.
 Among the 2,788 submissions reviewed, 1,788 were rejected and 1,000 were accepted.
-We prepared several attributes to be included for each question, answer, and hint, as shown in Tables 7, 8, and 9 in Appendix A.
-We discuss some of them in Section 4.
-Finally, we divided the questions and hints into train and test subsets.
-The train subset includes 4,500 hints for 900 questions, while the test subset contains 500 hints for 100 questions.
-Tables 10, 11, 12, and 13 in Appendix A show few examples of hints taken from the WikiHint dataset.
+3.2.3. Dataset Attributes
+We prepared several attributes to include for each question, answer, and hint, as shown in Tables 1, 2, and 3, respectively. We discuss some of them in Section 4.
+Two key attributes of a question are major and minor, which indicate the question type.
+To extract these attributes, we employ question classifiers, which categorize natural language questions into predefined classes based on their intended meaning.
+For example, the question Who is the director of the Godfather movie?
+falls into the PERSON:INDIVIDUAL class, while What is the capital of Austria?
+belongs to the LOCATION:CITY class.
+To develop an effective question classifier, we fine-tune the RoBERTa (Liu et al., 2019a) model using the TREC Question Classification dataset (Li and Roth, 2002), resulting in a model called QT Detector.
+This model achieves an accuracy of $92.8\%$ on the TREC Question Classification dataset [^5] The accuracy of the QT Detector on the TREC Question Classification dataset is $92.8\%$.
+Using QT Detector, questions are categorized into hierarchical classes with varying levels of granularity, encompassing 5 coarse-grained classes and 50 fine-grained classes.
+Finally, we divide the questions and hints into train and test subsets. The train subset includes 4,500 hints for 900 questions, while the test subset contains 500 hints for 100 questions.
+Figure 5. The HintRank method. First, the inputs are concatenated, and special tokens are added to format them for the BERT model. Finally, the BERT model determines which hint is more helpful.
+Figure 6. The distribution of the Train and Test subsets.
+4. Evaluation Approaches
+In this section, we first discuss the evaluation metrics we use to assess the quality of hints, followed by a detailed introduction to HintRank, our proposed automatic ranking method.
+4.1. Evaluation Metrics
+Mozafari et al.
+(2024b) introduced five evaluation metrics for assessing hint quality: Relevance, Readability, Convergence, Familiarity, and Answer Leakage.
+While the authors did not propose automatic methods for Relevance, Readability, and Answer Leakage, in this paper, we employ automated approaches for these metrics.
+In the following, we provide a detailed description of all metrics.
+Figure 7. Question difficulty based on the question types.
+Relevance
+This metric assesses the semantic connection between the hint and the question, ensuring that the hint remains meaningful and contextually appropriate.
+Hints can be considered a form of an answer since they provide explanations of the correct answer of the question.
+Based on this, one can evaluate the Relevance of a hint to its question as an Answer Relevance task (Es et al., 2024) - the task where the goal is to assess how pertinent the provided answer is to the target question.
+To compute the answer relevance metric, we employ DeepEval framework [^6] https://docs.confident-ai.com/ treating the hint as a kind of answer.
+Readability
+This metric measures the ease of understanding a hint or question, ensuring clarity and accessibility.
+It is crucial for making the hints effective without confusing users.
+To evaluate Readability (Liu and Lee, 2023), we finetune a RoBERTa (Liu et al., 2019b) model as a classifier on the OneStopEnglish dataset (Vajjala and Lučić, 2018).
+The finetuned model categorizes sentences into three classes: Beginner (0), Intermediate (1), and Advanced (2), reflecting their level of reading difficulty [^7] The accuracy of the readability estimator model is 62.3%.
+Convergence
+This metric measures how effectively a hint can narrow down or eliminate potential answers to the given question.
+To evaluate hints based on Convergence, we adopt the approach proposed by Mozafari et al.
+(2024b).
+In this method, candidate answers are first generated for a question.
+Then, the metric determines how many candidate answers are supported by the target hint.
+The convergence score is then derived on the basis of the capability of the hint to eliminate many incorrect candidate answers.
+We utilize two core models for convergence evaluation: LLaMA-3.1-8B and LLaMA-3.1-70B (Dubey et al., 2024).
+Table 4. Statistics of WikiHint dataset.
 Train Test
@@
 Avg. #entities / hint 1.2 1.18
-Table 1: Statistics of WikiHint dataset.
-Rank 1 2 3 4 5
-Average Length 16.99 17.67 18.02 18.14 18.3
-Table 2: Average length of hints vs. their ranks.
+Table 5. Question distribution by source and difficulty.
+Source ChatGPT Natural Questions SQuAD 2.0
+Difficulty 0.43 0.34 0.38
+Familiarity
+This metric measures the degree to which users are expected to know the information expressed in the hints.
+To evaluate hints based on Familiarity, we follow the approach proposed by Mozafari et al.
+(2024b).
+In this method, the named entities are first extracted from the evaluated hint, and the number of views for their corresponding Wikipedia pages is retrieved.
+After normalizing these values, the final familiarity score is computed as the average of the normalized view counts.
+Answer Leakage
+The last metric measures the degree to which a hint directly reveals the answer.
+This metric ensures that hints guide users toward the solution without explicitly disclosing it, promoting problem-solving rather than direct answer retrieval.
+To compute the degree of Answer Leakage, we assess the semantic similarity between each word in a hint and the corresponding answer using the RoBERTa model.
+The final score is obtained by maximizing and averaging these similarity values.
+4.2. HintRank Method
+In addition to the above automatic evaluation approaches involving individual hints, we introduce a new lightweight evaluation method, HintRank, for evaluating and ranking hints using pairwise preferences.
+Building on the success of widely-used automatic evaluation metrics like BERTScore (Zhang et al., 2020), BEM (Bulian et al., 2022), MoverScore (Zhao et al., 2019), and BLEURT (Sellam et al., 2020), which leverage BERT (Devlin et al., 2019) as the core evaluation module and demonstrate its effectiveness, we chose BERT as the base for the HintRank method.
+Our method determines the better hint within a pair of hints.
+Figure 5 illustrates the proposed method.
+In HintRank, we begin by concatenating a given question and its answer with two hints, labeled as $\text{Hint}_{1}$ and $\text{Hint}_{2}$ to create an input compatible with the BERT model.
+Note that in the answer-agnostic scenarios, we avoid appending the answer to the evaluated hints.
+Such constructed input is then processed by BERT model, which produces one of two possible outputs: 0 or 1.
+An output of 0 means that $\text{Hint}_{2}$ is of higher quality than $\text{Hint}_{1}$, whereas an output of 1 suggests that $\text{Hint}_{1}$ is superior to $\text{Hint}_{2}$.
+As HintRank operates on pairwise preferences, it requires $\binom{n}{2}$ comparisons for a question with $n$ hints, with a runtime complexity of $\mathcal{O}(n^{2})$.
+Table 6.
+Quality comparison of WikiHint and TriviaHG.
+Relevance, convergence, familiarity, and answer leakage are measured on a scale from 0 to 1, while readability is rated on a scale from 0 to 2 (the lower, the more readable).
 Dataset Subset Relevance Readability Convergence Familiarity Length
-Answer Leakage
-Degree (Avg)
-Answer Leakage
-Degree (Max)
+Answer Leakage Degree
+Answer Leakage Degree
 TriviaHG Entire 0.95 0.71 0.57 0.77 20.82 0.23 0.44
@@
 WikiHint Test 0.98 0.83 0.72 0.73 18.32 0.24 0.47
-Table 3: Quality comparison of WikiHint and TriviaHG.
-Relevance, convergence, familiarity, and answer leakage are measured on a scale from 0 to 1, while readability is rated on a scale from 0 to 2 (the lower, the more readable).
-4 Evaluation Approaches
-Mozafari et al.
-(2024) proposed several evaluation metrics for assessing the quality of hints, including Relevance, Readability, Convergence, and Familiarity, although the authors proposed automatic evaluation methods only for Convergence and Familiarity.
-Convergence is a measure of how effectively a hint can narrow down or eliminate potential answers to a given question.
-Familiarity measures the expected level of knowledge of information expressed in hints.
-To evaluate hints based on Convergence and Familiarity, we follow the method proposed by Mozafari et al.
-(2024).
-However, we employ two cores for convergence including LLaMA-3.1-8b and LLaMA-3.1-70b Dubey et al.
-(2024).
-We extend the above evaluation scheme by incorporating automatic methods for evaluating hint’s Relevance and Readability.
-We also propose a new metric for evaluating the probability of answer leakage - a case when a hint directly reveals the answer in its content.
-We introduce a lightweight automatic evaluation method for assessing hint quality in a pairwise scenario.
-The way to compute those additional metrics is briefly described below.
-Hints can be considered a form of an answer since they provide explanations of the question’s correct answer.
-Based on this, one can evaluate the Relevance of a hint to its question as an Answer Relevance task Es et al.
-(2024) - the task where the goal is to assess how pertinent the provided answer is to the target question.
-To compute the answer relevance metric, we employ DeepEval framework [^5] https://docs.confident-ai.com/ treating hint as a kind of answer.
-To evaluate Readability Liu and Lee (2023), we finetune a RoBERTa Liu et al.
-(2019) model as a classifier on the OneStopEnglish dataset Vajjala and Lučić (2018).
-The finetuned model categorizes sentences into three classes: Beginner (0), Intermediate (1), and Advanced (2), reflecting their level of reading difficulty [^6] The accuracy of the readability estimator model is 62.3%.
-To calculate the Answer Leakage Degree, we measure the semantic similarity between each word of a hint and an answer using RoBERTa model.
-4.1 HintRank Evaluation Approach
-In addition to the above automatic evaluation approaches involving individual hints, we introduce a new lightweight evaluation method, HintRank, for evaluating and ranking hints using pairwise preferences.
-Building on the success of widely-used automatic evaluation metrics like BERTScore Zhang et al.
-(2020), BEM Bulian et al.
-(2022), MoverScore Zhao et al.
-(2019), and BLEURT Sellam et al.
-(2020), which leverage BERT Devlin et al.
-(2019) as the core evaluation module and demonstrate its effectiveness, we chose BERT as the foundation for the HintRank method.
-Our method determines the better hint within a pair of hints.
-Figure 2 illustrates the proposed method.
-In the HintRank method, we begin by concatenating a given question and its answer with two hints, labeled as $\text{Hint}_{1}$ and $\text{Hint}_{2}$, to create an input compatible with BERT model.
-Note that in the answer-agnostic scenarios, we avoid appending the answer to the evaluated hints.
-Such constructed input is then processed by BERT model, which produces one of two possible outputs: 0 or 1.
-An output of 0 means that $\text{Hint}_{2}$ is of higher quality than $\text{Hint}_{1}$, whereas an output of 1 suggests that $\text{Hint}_{1}$ is superior to $\text{Hint}_{2}$.
-As HintRank operates on pairwise preferences, it requires $\binom{n}{2}$ comparisons for a question with $n$ hints, with a runtime complexity of $\mathcal{O}(n^{2})$.
-5 Experiments and Results
-5.1 Data Analysis
+5. Experiments and Results
+In this section, we first analyze the WikiHint dataset from several different perspectives.
+Following that, we present the results of the human evaluation.
+We then examine the performance of different LLMs on the dataset.
+Lastly, we compare the HintRank method and the convergence metric in different scenarios.
+5.1. Data Analysis
 The WikiHint dataset is split into a train set with 4,500 hints (900 questions) and a test set with 500 hints (100 questions).
-Table 1 provides the statistics of both train and test sets, while Figure 7 in Appendix A shows their distributions according to the question types, indicating that the distributions are well-matched.
-Figure 3: Average convergence of the hints of WikiHint based on the hint ranks.
+Table 4 provides the statistics of the train and test sets, while Figure 6 shows their distributions according to the question types, indicating that the distributions are well matched.
+Figure 8. Average convergence of the hints of WikiHint based on the hint ranks.
 We next analyze the difficulty levels of questions in WikiHint.
-To evaluate the difficulty, we utilize the Reference-based Question Complexity method Gabburo et al.
-(2024).
+To evaluate the difficulty, we utilize the Reference-based Question Complexity method (Gabburo et al., 2024).
 This method computes the difficulty of a question by assessing how many of its retrieved passages contain the correct answer and by measuring the relevance between the retrieved passages and the question.
 It then calculates the difficulty score for the question based on such computed features.
-In particular, we use the DPR method Karpukhin et al.
-(2020) as the retrieval technique, employing an English Wikipedia dump preprocessed by Karpukhin et al.
+We use the DPR method (Karpukhin et al., 2020) as the retrieval technique, employing the English Wikipedia dump preprocessed by Karpukhin et al.
 (2020) as the evidence source, and consider the top 30 most relevant passages as the retrieved passages.
-Figure 8 in Appendix A illustrates the computed question difficulty of WikiHint for train and test subsets [^7] We classify questions with difficulty scores below 0.33 as easy, those above 0.66 as hard and the rest as medium..
+Figure 7 illustrates the computed question difficulty of WikiHint for the train and test subsets [^8] We classify questions with difficulty scores below 0.33 as easy, those above 0.66 as hard and the rest as medium..
 The figure indicates that medium-hard questions are the most common as well as the train and test subsets have quite similar distributions in terms of question difficulty.
-Also, Table 6 in Appendix A highlights the difficulty levels of questions generated or extracted from various sources.
+Also, Table 5 highlights the difficulty levels of questions generated or extracted from various sources.
+Table 7. Demographic information of evaluators
+Evaluator Gender Age Education Level
+1 Female 18 High School
+2 Female 36 High School
+3 Female 30 Bachelor’s Degree
+4 Male 40 Master’s Degree
+5 Male 33 PhD
+Table 8. Average length of hints vs. their ranks.
+Rank 1 2 3 4 5
+Average Length 16.99 17.67 18.02 18.14 18.3
 Table 8 reveals an interesting insight regarding the length of hints, which can be considered as one of indicators of helpfulness.
-The results suggest that high-quality hints tend to be shorter in length (measured by the number of words) than the lower quality hints.
-This finding indicates an inverse correlation between hint length and helpfulness, challenging the intuition that longer hints are more informative or specific, and therefore more useful.
+The results suggest that high-quality hints tend to be shorter in length (measured by the number of words) than the lower-quality hints.
+This finding indicates an inverse correlation between the hint length and helpfulness, challenging the intuition that longer hints are more informative or specific, and therefore more useful.
 In contrast, shorter hints appear to be more concise and easier to follow, likely presenting more helpful information in the first place.
-We also evaluate the hints in the entire WikiHint dataset (and separately, in its train and test subsets) using the relevance, readability, convergence, familiarity, length, and answer leakage degree.
-We then compare these values with the ones obtained for TriviaHG dataset Mozafari et al.
-(2024) - the only existing hint dataset.
-The comparison results are presented in Table 3.
-The results indicate that in terms of relevance, readability, answer leakage degree, and familiarity, the metrics are nearly same between the two datasets.
-However, WikiHint has better convergence values compared to TriviaHG.
-Additionally, the hints in WikiHint are shorter in length, as measured by word count, than TriviaHG.
-These results indicate that the hints in WikiHint are of higher quality.
-Lastly, Figure 3 demonstrates the negative correlation between the convergence scores of hints and their helpfulness as represented by hint ranks assigned by crowdworkers.
-The plot suggests that the convergence scores can be considered a reliable metric for evaluating the helpfulness of hints and for hint ranking.
-5.2 Human Evaluation
-To manually evaluate hints, we recruited five independent evaluators, who were not involved in the dataset generation process, to answer the questions from the test subset of the WikiHint.
-The process was as follows: 1.
-Participants were asked to answer the question without using any hints.
-If they provided a correct answer, they proceeded to the next question.
-2.
-If they could not answer the question correctly, they were asked to review the hints until they could find the correct answer.
-By providing the correct answer, the participants could move to the next question.
-3.
-If the participants could not answer the question after reviewing all the hints, they were allowed to skip the question.
-Figure 4: The results of human evaluation.
+Table 9. Evaluation of generated hints based on relevance, readability, convergence, familiarity, length, and answer leakage across different scenarios.
 Model Config
@@
 LLaMA-3.1-8b Vanilla 0.78 1.76 0.04 0.3 0.83 52.99 0.22 0.5
-Table 4: Evaluation of generated hints based on relevance, readability, convergence, familiarity, length, and answer leakage across different scenarios.
-Figure 4 illustrates that all the participants could answer more questions across all of question types such as HUMAN, ENTITY, and LOCATION [^8] We use names as stated in the original dataset.
+We also evaluate the hints in the entire WikiHint dataset (and separately, in its train and test subsets) using the relevance, readability, convergence, familiarity, length, and answer leakage degree.
+We then compare these values with the ones obtained for the TriviaHG dataset (Mozafari et al., 2024b) - the highest-quality and largest hint dataset available.
+The comparison results are presented in Table 6.
+The results indicate that in terms of relevance, readability, answer leakage degree, and familiarity, the metrics are nearly the same between the two datasets.
+However, WikiHint has better convergence values compared to TriviaHG.
+Additionally, the hints in WikiHint are shorter, as measured by the word count.
+These results indicate that the hints in WikiHint are of higher quality.
+Lastly, Figure 8 demonstrates the correlation between the convergence scores of hints and their helpfulness as represented by hint ranks assigned by crowdworkers.
+The plot suggests that the convergence scores can be considered a reliable metric for evaluating the helpfulness of hints and can be used for hint ranking.
+In other words, the higher the convergence, the higher the rank of the hint.
+5.2. Human Evaluation
+To manually evaluate hints, we recruited five independent evaluators, whose backgrounds are shown in Table 7.
+The evaluators were not involved in the dataset generation process and were tasked with answering questions from the test subset of WikiHint, which consists of 100 questions and 500 hints.
+The evaluation process followed these steps:
+  1. (1)
+    Participants were asked to answer the question without using any hints. If they provided a correct answer, they proceeded to the next question.
+  2. (2)
+    If they could not answer the question correctly, they were asked to review the hints until they could find the correct answer.
+By providing the correct answer, the participants could move to the next question.
+  3. (3)
+    If the participants could not answer the question after reviewing all the hints, they were allowed to skip the question.
+Figure 9. The results of human evaluation.
+Figure 9 illustrates that all the participants could answer more questions in question types such as HUMAN, ENTITY, and LOCATION [^9] We use names as stated in the original dataset.
 when they used hints compared to the case without hints.
@@
 This suggests that generating effective hints becomes progressively more challenging for human, entity, and location questions, in that order.
-5.3 Model Performance
-To further assess the quality of hints, we analyze how well LLMs can automatically generate hints for questions.
-We use the open-source LLaMA models: LLaMA-3.1-8b, LLaMA-3.1-70b, and LLaMA-3.1-405b Dubey et al.
-(2024), and GPT-4 OpenAI et al.
-(2023) as the most powerful closed-source LLM for comparison.
-To explore different scenarios, we finetune [^9] We perform model finetuning using the API functions available on together.ai LLaMA-3.1-8b and LLaMA-3.1-70b on the training subset of the WikiHint dataset to evaluate the LLMs’ capabilities in hint generation when trained specifically on this task.
+Figure 10. Average length of hints generated by LLMs.
+5.3. Model Performance
+To further assess the quality of hints in our dataset, we analyze how well LLMs can automatically generate hints for input questions.
+We use the open-source LLaMA models: LLaMA-3.1-8b, LLaMA-3.1-70b, and LLaMA-3.1-405b (Dubey et al., 2024), and GPT-4 (OpenAI et al., 2023).
+Figure 11. Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
+To explore different scenarios, we finetune [^10] We perform model finetuning using the API functions available on together.ai.
+LLaMA-3.1-8b and LLaMA-3.1-70b on the train subset of the WikiHint dataset to evaluate the LLMs’ capabilities in hint generation when trained specifically on this task.
 For each question, we assign a hint as the target during the finetuning process.
 As a result of this learning strategy, during the inference stage, the finetuned model is prompted to generate one hint for each question.
-We consider two finetuning approaches: answer-aware and answer-agnostic.
-Given that LLMs typically handle most knowledge questions correctly, the answer-agnostic approach might be sufficient for generating hints.
-Besides, users generally do not know the answers to their questions when seeking hints.
-However, the answer-aware approach has its own advantages, too, such as in educational contexts where a teacher might use it to collect materials for class preparation.
-Due to the importance of both approaches, we chose to investigate fine-tuning of the LLMs in these two distinct scenarios.
-We found that shorter prompts were more effective in achieving the desired task.
-Longer, more detailed instructions often led to the model disregarding the key goal, i.e., generating hints, and instead focusing on irrelevant details.
-In contrast, shorter prompts increased the likelihood of successful task completion.
-After experimenting, we opted for the following prompt as the system prompt:
-You are a hint generator for the factoid questions. The user asks you a question and you should generate a hint for that question without revealing the answer in the hint.
-Two distinct user prompts were employed to generate hints within a zero-shot learning strategy.
-Assuming a question $q$ as an input, the answer-agnostic prompt was ‘Give me the best hint for this question: $q$.
-The answer-aware prompt included the answer $a$ as follows: ‘Give me the best hint for this question: $q$?
-The answer for the question is $a$’.
-Figure 5: Average length of generated hints by LLMs.
-To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA where wA means With Answer and woA means Without Answer.
-We test these models on the WikiHint test to assess the impact of finetuning and the inclusion of answers in the prompt.
-Figure 5 illustrates that as LLMs decrease in size and hint generation capability, the length of the generated hints increases.
-This supports our observation made in Section 5.1 of an inverse correlation between hint length and hint quality.
-Additionally, hints produced by finetuned models are generally shorter than those from vanilla models, indicating that finetuned models may generate higher-quality hints.
-Moreover, hints in the answer-aware scenarios are shorter compared to those in answer-agnostic scenarios, suggesting that when the answer is provided along with the question, LLMs are able to produce more effective hints.
-Table 4 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity.
-The results indicate that more powerful LLMs are capable of generating more relevant hints.
-Regarding readability, GPT-4 exhibits the highest quality, followed closely by LLaMA-3.1-405b and LLaMA-3.1-70b, while LLaMA-3.1-8b shows the lowest readability.
-It also demonstrates that more powerful LLMs can generate more readable hints.
-Additionally, finetuned models consistently outperform their vanilla counterparts, and answer-aware prompts yield better results compared to answer-agnostic prompts for readability and familiarity.
-The answer leakage degree indicates that the prompt we use is effective in preventing LLMs from including answers, their synonyms or very similar terms in the generated hints, as the results closely align with those of WikiHint shown in Table 3.
-Figure 6 illustrates that as LLMs decrease in size and capability, the convergence of their hints also diminishes.
-This trend is observed for both LLaMA-3.1-8b and LLaMA-3.1-70b, used as the cores of the convergence method.
-This supports our claim in Section 5.1 regarding the correlation between convergence and ranks.
-The figure also shows that the average convergence for the answer-aware approach surpasses that of the answer-agnostic approach, suggesting that including the answer in the prompt makes it easier for LLMs to generate hints.
-Furthermore, LLMs finetuned on the train subset of the WikiHint dataset achieve better convergence scores than their vanilla counterparts, indicating the efficacy of WikiHint for finetuning LLMs in hint generation.
-Figure 6: Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
-5.4 HintRank Evaluation Method
+Table 10. Comparison between Convergence metric, LLM-based ranking, and HintRank.
 Method Config Use Answer Accuracy (%) Correlation (%)
@@
 HintRank FTwA 68.55 52.34
-Table 5: Comparison between Convergence metric, LLM-based ranking, and HintRank.
+We consider two finetuning approaches depending on whether the answer is or is not used in the input to LLMs: Answer-Aware and Answer-Agnostic.
+Given that LLMs typically handle most knowledge questions correctly, the answer-agnostic approach might be sufficient for generating hints.
+Besides, users generally do not know the answers to the questions for which they need hints.
+However, the answer-aware approach still has its own use, such as in educational contexts where a teacher might wish to collect materials for class preparation.
+Due to the importance of both approaches, we chose to investigate fine-tuning of the LLMs in these two distinct scenarios.
+We found that shorter prompts were more effective in achieving the desired task.
+Longer, more detailed instructions often led to the model disregarding the key goal of generating hints, and instead focusing on irrelevant details.
+In contrast, shorter prompts increased the likelihood of successful task completion.
+After experimenting, we opted for the following prompt as the system prompt:
+You are a hint generator for the factoid questions. The user asks you a question and you should generate a hint for that question without revealing the answer in the hint.
+Two distinct user prompts were employed to generate hints within a zero-shot learning strategy.
+Assuming a question $q$ as an input, the answer-agnostic prompt was ‘Give me the best hint for this question: $q$.
+The answer-aware prompt included the answer $a$ as follows: ‘Give me the best hint for this question: $q$?
+The answer to the question is $a$’.
+To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA, where FT stands for Finetuned, wA denotes With Answer, and woA means Without Answer.
+We test these models on the WikiHint test set to analyze the impact of finetuning, the inclusion of answers in the prompt, and their overall effectiveness in generating informative hints.
+Figure 10 illustrates that as LLMs increase in their size and hint generation capability, the length of the generated hints decreases.
+This supports our observation made in Section 5.1 of an inverse correlation between hint length and hint quality.
+Additionally, hints produced by finetuned models are generally shorter than those from vanilla models, indicating that finetuned models may generate higher-quality hints.
+Moreover, hints in the answer-aware scenarios are shorter compared to those in answer-agnostic scenarios, suggesting that when the answer is provided along with the question, LLMs are able to produce more effective hints.
+Figure 12.
+Accuracy of HintRank for different hint pairs in different scenarios.
+The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.
+Table 9 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity.
+The results indicate that more powerful LLMs are capable of generating more relevant hints.
+Regarding readability, GPT-4 exhibits the highest quality, followed closely by LLaMA-3.1-405b and LLaMA-3.1-70b, while LLaMA-3.1-8b shows the lowest readability.
+It also demonstrates that more powerful LLMs can generate more readable hints.
+Additionally, finetuned models consistently outperform their vanilla counterparts, and answer-aware prompts yield better results compared to answer-agnostic prompts for readability and familiarity.
+The answer leakage degree indicates that the prompt we use is effective in preventing LLMs from directly including answers, their synonyms or very similar terms in the generated hints, as the results closely align with those of WikiHint shown in Table 6.
+Figure 11 illustrates that as LLMs increase in size and capability, the convergence of their hints also increases.
+This trend is observed for both LLaMA-3.1-8b and LLaMA-3.1-70b, used as the cores of the convergence method.
+This supports our claim in Section 5.1 regarding the correlation between convergence and ranks.
+The figure also shows that the average convergence of hints created by the answer-aware approach surpasses that of the ones by answer-agnostic approach, suggesting that including the answer in the prompt makes it easier for LLMs to generate good hints.
+Furthermore, LLMs finetuned on the train subset of the WikiHint dataset achieve better convergence scores than their vanilla counterparts, indicating the usefulness of WikiHint for finetuning LLMs for hint generation.
+5.4. Evaluation of HintRank Method
 As outlined in Section 4, we also propose in this paper a novel evaluation method, HintRank, for ranking hints using the BERT model.
-Alongside finetuning BERT, we additionally finetune LLaMA-3.1-8b and LLaMA-3.1-70b models on the train set of the WikiHint to assess the performance of these LLMs in identifying high-quality hints.
-Similar to the experiments described in Section 5.3, we examine various scenarios including answer-aware and answer-agnostic contexts, and compare vanilla models with their finetuned counterparts.
+In addition to finetuning BERT, we additionally finetune LLaMA-3.1-8b and LLaMA-3.1-70b models on the train set of the WikiHint to assess the performance of these LLMs in identifying high-quality hints.
+Similarly to the experiments described in Section 5.3, we examine various scenarios including answer-aware and answer-agnostic contexts, and compare vanilla models with their finetuned counterparts.
 We use the following prompt as the system prompt:
-You are a hint evaluator for the factoid questions. The user gives you a question and two hints and you should specify which hint for that question is a better hint and more helpful.
-Two distinct user prompts are employed to evaluate hints within a zero-shot learning strategy.
+You are an expert hint evaluator for factoid questions. Given a question and two hints, your task is to determine which hint is more helpful in guiding the user toward the correct answer.
+Two distinct user prompts are employed for evaluating hints within a zero-shot learning strategy.
 Assuming a question $q$ as a question and $h_{1}$ and $h_{2}$ as a pair of hints, the answer-agnostic prompt is:
-Which hint is better to find the answer of this question: $q$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
+Which hint is better to find the answer of this question: $q$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between ”Hint_1” and ”Hint_2” without any explanations.
 and the answer-aware prompt with answer $a$ is:
-Which hint is better to find the answer of this question: $q$.
-The answer for this question is $a$.
-Hint_1: $h_{1}$.
-Hint_2: $h_{2}$.
-Just choose between "Hint_1" and "Hint_2" without any explanations.
-We benchmark HintRank against the Convergence metric which turned out to be useful for hint ranking assessment as indicated in Figure 3.
-To convert pairwise rankings to listwise rankings, we apply the Bradley–Terry model Bradley and Terry (1952).
-We evaluate the correlation between the rankings with Pearson Correlation Mining (2006).
-Table 5 outlines the key features and differences among various scenarios.
+Which hint is better to find the answer of this question: $q$. The answer to this question is $a$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between ”Hint_1” and ”Hint_2” without any explanations.
+We benchmark HintRank against the Convergence metric which turned out to be useful for evaluating hint ranking as indicated in Figure 8.
+To convert pairwise rankings to listwise rankings, we apply the Bradley–Terry model (Bradley and Terry, 1952).
+We then evaluate the correlation between the rankings with the Pearson Correlation (Mining, 2006).
+Table 10 outlines the key features and differences between various scenarios.
 The results indicate that with the increase in the size and power of LLMs, both accuracy and correlation improve.
@@
 Moreover, finetuned versions outperform their vanilla counterparts, demonstrating that the WikiHint dataset is well-suited for model fine-tuning to rank hints.
-Surprisingly, the BERT-base method outperforms LLMs, including finetuned versions in the answer-aware scenario.
+Surprisingly, the BERT-base method outperforms LLMs, including their finetuned versions in the answer-aware scenario.
 This holds true for both Bert-FTwoA and Bert-FTwA, although BERT-base also performs better in the answer-aware approach compared to answer-agnostic.
@@
 Utilizing BERT-based models instead of LLMs enhances the speed and accessibility of HintRank, reducing computational demands.
-Figure 13 in Appendix A shows that accuracy improves as the rank difference between hints increases, indicating it’s harder to correctly order hints with closer ranks.
-6 Conclusions
-In this paper, we introduced the first manually created dataset for hint generation and hint ranking.
-We also presented a new lightweight method for evaluating and ranking hints.
+Figure 12 shows that the accuracy improves as the rank difference between the hints increases, indicating it is harder to correctly order hints with closer ranks.
+6. Conclusions
+Automatic hint generation is a task closely related to question answering and question generation problems.
+Effective hints can constitute a useful mechanism for engaging users in the information seeking process.
+In this paper, we introduced the first manually created dataset for hint generation and ranking.
+We also presented a new lightweight method for ranking and evaluating hints.
 To demonstrate the effectiveness of our dataset, we conducted experiments where humans attempted to answer questions with and without the use of hints.
@@
 We then finetuned LLMs using our dataset, prompting them to generate new hints for different questions.
-The high quality of the generated hints indicates that our dataset is well-suited for finetuning LLMs for HG task.
-We also finetuned BERT and LLMs on the dataset for the task of hint ranking and evaluated their performance.
+The high quality of the generated hints indicates that our dataset is well-suited for finetuning LLMs for the HG task.
+We also finetuned BERT and LLMs on WikiHint for the task of hint ranking and evaluated their performance.
 The results reveal that encoder-based models outperform decoder-based models in hint ranking.
-In future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.
-Limitations
-Our study has the following limitations:
-  • The need for generative capabilities in hint generation task necessitates the use of LLMs.
-However, this dependency is a limitation as fine-tuning and prompting LLMs require extensive computational resources and are time-consuming.
-  • Our research focus on factoid questions may limit its applicability to other types of questions that involve more complex or abstract answers.
-Factoid questions, by their nature, provide clear and concrete answers, which simplifies automated hint generation and evaluation but may not fully capture the breadth of human inquiry.
-  • The WikiHint dataset is exclusively written in the English language.
-While this facilitates accessibility for a global audience and ensures compatibility with most existing Large Language Models, it also limits the dataset’s applicability in multilingual or non-English contexts, potentially excluding non-English speakers and diverse linguistic data.
-Ethical Considerations
-Our study utilizes GPT models, which are covered by the OpenAI License and Apache-2.0 license, and the LLaMA model, which is distributed under Meta’s LLaMA 2 Community License Agreement.
-We comply with these licensing agreements in all applications.
-Additionally, the datasets we use are sourced from repositories that are approved for academic use.
-The artifacts developed during our research are made available under the MIT license to facilitate straightforward modifications and use by the research community.
-We ensure that our data management, model training, and dissemination practices meet ethical standards and legal requirements associated with each artifact we use.
-Appendix A Appendix
-Figure 7: The distribution of Train and Test subsets.
-Figure 8: Question difficulty based on different question types.
-Source ChatGPT NQ SQuAD 2.0
-Difficulty 0.43 0.34 0.38
-Table 6: Question distributions based on the their sources and difficulty.
-Figure 9: The MTurk Worker interface for the hint generation task.
-Figure 10: The summarized instructions for the hint generation task.
-Figure 11: The detailed instructions for the hint generation task.
-Figure 12: Good and bad examples for the hint generation task.
-Attribute Description
-question The content of the question.
-major The major category of the question.
-minor The specific sub-category of the question.
-entity Content of an entity
-ent_type Type of entity (e.g., GPE, PERSON).
-start_index Start index of the entity in the question.
-end_index End index of the entity in the question.
-wikipedia_page_title The title of the corresponding Wikipedia page to the entity in question.
-wiki_views_per_month Number of views per month of the Wikipedia page for the entity in question.
-normalized_views Views normalized to scale from 0 to 1 for the entity in the question.
-readability Indicates the readability score
-familiarity Indicates the familiarity score
-difficulty Indicates the difficulty level of the question
-Table 7: A detailed description of attributes of a question in WikiHint.
-Attribute Description
-answer The actual answer.
-entity Content of an entity as identified within the answer.
-ent_type Type of entity.
-start_index Start index of the entity in the answer.
-end_index End index of the entity in the answer.
-wikipedia_page_title The title of the corresponding Wikipedia page to the entity in the answer.
-wiki_views_per_month Number of views per month of the Wikipedia page for the entity in the answer.
-normalized_views Views normalized to scale from 0 to 1 for the entity in the answer.
-familiarity Indicates the familiarity score
-difficulty Indicates the difficulty level of the answer
-Table 8: A detailed description of attributes of an answer in WikiHint.
-Attribute Description
-hint Hint provided to assist with the question.
-source URL source of the hint.
-entity Content of entities mentioned in the hint.
-ent_type Category of each entity (e.g., PERSON, GPE) mentioned in the hint.
-start_index Specific start index where the entity is found in the hint text.
-end_index Specific end index where the entity is found in the hint text.
-wikipedia_page_title The title of the corresponding Wikipedia page to the entities in the hint.
-wiki_views_per_month Number of views per month of the Wikipedia pages for the entities of the hint.
-normalized_views Views normalized to scale from 0 to 1 for the entity in the hint.
-relevance Indicates the relevance score
-readability Indicates the readability score
-convergence Indicates the convergence score
-familiarity Indicates the familiarity score
-answer_leakage Indicates the answer leakage score
-rank Priority or helpfulness rating of the hint
-Table 9: A detailed description of attributes of a hint in WikiHint.
-Question What is the driest country in sub-Saharan Africa?
-Fifth Best Hint The earliest settlers in this country were in the 18th century and crossed the Orange River to move into the area.
-Fourth Best Hint The name of this country is derived from the oldest desert on Earth.
-Third Best Hint This country’s coat of arms is a shield with the same design as the flag, 2 antelopes and a red blue and white bird.
-Second Best Hint This country is a country in southern Africa with its western border along the Atlantic Ocean.
-Best Hint This country’s flag is blue and green with a red stripe down the middle and the symbol of a sun in the top left corner.
-Table 10: First example of the WikiHint.
-Question What artist received the Presidential Medal of Freedom posthumously in 2018, in recognition of his contributions to American culture?
-Fifth Best Hint The artist had karate as a lifelong interest, including its moves in his performances.
-Fourth Best Hint The artist served in the Army after he became famous.
-Third Best Hint The artist never performed outside North America.
-Second Best Hint The artist tried to establish a career in films with pictures such as Jailhouse Rock and Fun in Acapulco.
-Best Hint The artist is known for hits such as "Love Me Tender" and "Jailhouse Rock".
-Table 11: Second example of the WikiHint.
-Question What British comedy team is famous for its ’Four Yorkshiremen’ sketch, portraying exaggerated tales of hardship with humor?
-Fifth Best Hint A giant cupid’s foot is repeatedly used in the show.
-Fourth Best Hint This British comedy troupe is formed in 1969 consisting of 6 members.
-Third Best Hint The team was awarded the AFI Star Award by the American Film Institute in 1998.
-Second Best Hint The team received the BAFTA Award for Outstanding British Contribution to Cinema at the 41st British Academy Film Awards in 1988.
-Best Hint The Holy Grail and Life of Brian are some of their greatest comedy films.
-Table 12: Third example of the WikiHint.
-Question What fictional pirate is known for carrying a compass that doesn’t point north but rather to what the user wants most?
-Fifth Best Hint One of the pirate’s most famous quotes is "Now, bring me that horizon".
-Fourth Best Hint The fictional pirate appeared in the video game series "Kingdom Hearts".
-Third Best Hint The fictional pirate has a blood debt to another character called Davy Jones.
-Second Best Hint The ship of the fictional pirate is called the Black Pearl.
-Best Hint The fictional pirate is played by Johnny Depp in the film series.
-Table 13: Fourth example of the WikiHint.
-A hint must not include the exact answer explicitly.
-A hint must be a sentence.
-A hint must be specific, not generic.
-A hint must be from the corresponding Wikipedia page.
-A hint must have a unique rank.
-Table 14: A detailed criteria or standards used for verifying the generated hints during the selection process.
-Figure 13: Accuracy of HintRank for different hint pairs in different scenarios.
-The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.
+In the future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.

### v3

@@
 We first introduce a manually constructed hint dataset, WikiHint, which is based on Wikipedia and includes 5,000 hints created for 1,000 questions.
-We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts.
+We then finetune open-source LLMs for hint generation in answer-aware and answer-agnostic contexts.
 We assess the effectiveness of the hints with human participants who answer questions with and without the aid of hints.
 Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings.
-Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves quality of generated hints, and (c) encoder-based models perform better than decoder-based models in hint ranking.
-Hint Dataset, Hint Evaluation, Hint Generation
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
+Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves the quality of generated hints, and (c) encoder-based models perform better than decoder-based models in hint ranking.
+Hint Dataset, Hint Evaluation, Hint Generation, Hint Ranking
+journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730284isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
 1. Introduction
@@
 Hints for a question from the WikiHint dataset, with their corresponding rankings (1 being the highest and 5 the lowest) which let users find the answer.
-Note that the opposite arrangement of the hints (from 1 to 5) would make the answer finding task easiest.
+Note that the opposite arrangement of the hints would make the answer finding task easiest when hints are read from rank 1 to 5.
+[…]
 Figure 2. Pipeline of WikiHint dataset generation. The numbers in the arrows indicate the counts of output questions. Qs and Hs denote the number of questions and hints, respectively.
+[…]
 In recent years, question answering (QA) systems have risen in importance, offering users the opportunity to effortlessly receive answers to different types of questions (Mavi et al., 2024; Karpukhin et al., 2020; Zhao et al., 2021; Abdel-Nabi et al., 2023).
@@
 For this purpose, we select Wikipedia as the primary source, as it is one of the most widely recognized and comprehensive repositories of factual information.
-In this paper, we propose the first manually curated dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint, which has been constructed and designed for hint generation, ranking and evaluation.
+In this paper, we propose the first manually curated dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint, which has been constructed and designed for hint generation, ranking, and evaluation.
 Figure 1 shows a question from the WikiHint dataset, with its hints and the corresponding ranks reflecting the usefulness of each hint.
@@
 \${question} represents the question presented to the worker, \${answer} is the corresponding answer, and \${link} provides the link to the relevant Wikipedia page.
+[…]
 3.1. Question Sampling Module
@@
 Figure 4. The instructions for the hint generation and ranking tasks on the Amazon MTurk platform.
+[…]
 Table 1. A detailed description of attributes of a question.
@@
 normalized_views Views normalized to scale from 0 to 1 for the entity in the question.
-readability The readability score.
-familiarity The familiarity score.
+readability The readability score of the question.
+familiarity The familiarity score of the question.
 difficulty The difficulty level of the question.
@@
 normalized_views Views normalized to scale from 0 to 1 for the entity in the answer.
-familiarity The familiarity score.
-difficulty The difficulty level of the answer.
+familiarity The familiarity score of the answer.
+[…]
 Table 3. A detailed description of attributes of a hint.
@@
 rank Priority or helpfulness rating of the hint.
+[…]
 3.2.2. Hint Verification
@@
 Figure 5. The HintRank method. First, the inputs are concatenated, and special tokens are added to format them for the BERT model. Finally, the BERT model determines which hint is more helpful.
+[…]
 Figure 6. The distribution of the Train and Test subsets.
+[…]
 4. Evaluation Approaches
-In this section, we first discuss the evaluation metrics we use to assess the quality of hints, followed by a detailed introduction to HintRank, our proposed automatic ranking method.
+In this section, we begin by introducing the evaluation metrics used to assess the quality of hints.
+These metrics are available through the HintEval (Mozafari et al., 2025) framework, which provides a standardized setup for hint evaluation [^6] https://github.com/DataScienceUIBK/HintEval.
+Following this, we present HintRank, our proposed method for automatic hint ranking.
 4.1. Evaluation Metrics
@@
 Figure 7. Question difficulty based on the question types.
+[…]
 Relevance
@@
 We utilize two core models for convergence evaluation: LLaMA-3.1-8B and LLaMA-3.1-70B (Dubey et al., 2024).
+Familiarity
+This metric measures the degree to which users are expected to know the information expressed in the hints.
+To evaluate hints based on Familiarity, we follow the approach proposed by Mozafari et al.
+(2024b).
+In this method, the named entities are first extracted from the evaluated hint, and the number of views for their corresponding Wikipedia pages is retrieved.
+After normalizing these values, the final familiarity score is computed as the average of the normalized view counts.
 Table 4. Statistics of WikiHint dataset.
@@
 Difficulty 0.43 0.34 0.38
-Familiarity
-This metric measures the degree to which users are expected to know the information expressed in the hints.
-To evaluate hints based on Familiarity, we follow the approach proposed by Mozafari et al.
-(2024b).
-In this method, the named entities are first extracted from the evaluated hint, and the number of views for their corresponding Wikipedia pages is retrieved.
-After normalizing these values, the final familiarity score is computed as the average of the normalized view counts.
 Answer Leakage
@@
 Figure 8. Average convergence of the hints of WikiHint based on the hint ranks.
+[…]
 We next analyze the difficulty levels of questions in WikiHint.
@@
 Figure 9. The results of human evaluation.
+[…]
 Figure 9 illustrates that all the participants could answer more questions in question types such as HUMAN, ENTITY, and LOCATION [^10] We use names as stated in the original dataset.
@@
 Figure 10. Average length of hints generated by LLMs.
+[…]
 5.3. Model Performance
@@
 Figure 11. Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
+[…]
 To explore different scenarios, we finetune [^11] We perform model finetuning using the API functions available on together.ai.
@@
 The answer to the question is $a$’.
-To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA, where FT stands for Finetuned, wA denotes With Answer, and woA means Without Answer.
+To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches including Vanilla-wA, Vanilla-woA, FTwA, and FTwoA, where FT stands for Finetuned, wA denotes With Answer, and woA means Without Answer.
 We test these models on the WikiHint test set to analyze the impact of finetuning, the inclusion of answers in the prompt, and their overall effectiveness in generating informative hints.
@@
 The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.
+[…]
 Table 9 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity.
@@
 Figure 12 shows that the accuracy improves as the rank difference between the hints increases, indicating it is harder to correctly order hints with closer ranks.
-6. Conclusions
+6. Conclusion
 Automatic hint generation is a task closely related to question answering and question generation problems.
@@
 The results reveal that encoder-based models outperform decoder-based models in hint ranking.
-In the future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.
+In the future, we plan to generate personalized hints tailored to the knowledge of users. The main challenge here will be to develop appropriate datasets and solutions for user profiling.